# Flexible Music-Conditioned Dance Generation with Style Description Prompts

### v1

@@ V-E Results of Flexible Dance Generation
V-F Discussions
We conduct detailed ablation studies, analyze the impacts of audio representations and various style embeddings, and discuss the limitations of evaluation metrics.
Audio Representation: We evaluate three independent feature extraction strategies: Jukebox [47], Encodec [48] and Librosa [49].
Jukebox [47], a pre-trained generative model designed for music generation, showcasing notable performance in tasks specific to music prediction as evidenced in prior studies [6]. Musicians and composers can use Jukebox to create original musical compositions, exploring and innovating between different genres and styles. Inspired by these advances, here, we extract Jukebox music features as music conditions.
Encodec [48], an advanced real-time, high-fidelity, audio codec. Encodec is a pre-trained model, which has been used as music conditional feature extractor in dance generation tasks as well. It achieves state-of-the-art quality levels in both voice and music compression at multiple audio compression ratios and sample rates.
Librosa [49] is a library for analyzing audio signals. It provides a rich set of functions including loading audio files, extracting features, computing rhythmic features and performing audio signal processing. Here, we use librosa to extract the following musical features: envelope (1-dim), MFCC (20-dim), chroma (12-dim), one-hot peaks (1-dim) and one-hot beats (1-dim).
[…]
It is demonstrated through ablation experiments that our method achieves the best Beat Alignment Score and maintains superior performance in other metrics.
@@ V-F Discussions
Visualizations: For the same music, our approach can generate a variety of sensible dance movements. Some of the dance visualization results are shown in Figure 4. Here, we list generated dances for four types of music: Breaking, Krump, Locking and Waacking. For each type, we list five dance movements. As can be seen from the figure, our approach can generate diverse and plausible dance movements.
The beat alignment score calculates the average time distance between each music beat and its closest dance beat. However, dancing is not solely about syncing local speed minima in joints with beats. Instead, musical beats serve as a flexible reference for timing, rhythm, and the smooth transition between dance movements and steps. While this metric has driven progress on this issue in the past, more dance-specific metrics are needed to realize breakthroughs in this area.
VI Conclusion
@@
In this work, we study music-conditioned dance generation and introduce a transformer-based diffusion framework called Dance Generation with Style Description Prompts (DGSDP) which can generate arbitrary dance sequences. The DGSDP mainly comprises Music-Conditioned Style-Aware Diffusion (MCSAD) and Style Modulation with Description Prompts. Our approach obtains state-of-the-art performance on a variety of tasks such as dance generation, long-term dance generation, dance in-betweening, dance inpainting. Ablated experiments demonstrate the effectiveness of the modulation module and the style description prompts. We also conduct a thorough investigation into various encodings of both textual music style and the audio music. To obtain dances that align better with the music, we suggest using Jukebox to extract audio representations and applying style description prompts augmented by large language models. The proposed framework is suitable for flexible dance generation, enabling the user to freely edit or generate the desired dance sequence. We hope that this work will inspire further research in the fields of automatic dance generation and interactive dance generation.

### v2

@@ IV-G Ablation Studies and Discussions
Figure 6: Visualization of generated dances for the same piece of music with different style prompts. In each example, we present dances in five different styles.
User Study: In addition to the quantitative experiments mentioned above, we invite human users to assess the visual quality of dance movements created by our method and by EDGE [9]. These users come from a diverse range of regions, ages, and genders. When presented with dance movements generated from the same music clip, the users are asked: ”Which dance do you prefer?” Furthermore, we ask distinct users to articulate their specific feelings. Our findings reveal that 60% of the users prefer dance movements generated by our method compared to those of EDGE. Additionally, the movements created by our method exhibit greater diversity, and even some samples surpass the ground truth in quality.
@@
Visualizations: For the same music, our approach can generate a variety of sensible dance movements. Some of the dance visualization results are shown in Figure 5. Here, we list generated dances for two types of music: Breaking and Locking. For each type, we list five dance movements. As can be seen from the figure, our approach can generate diverse and plausible dance movements.
We also provide qualitative results of generated dances of arbitrary styles Figure 6. Our method is capable of generating more diverse and expressive dances, holding considerable potential for creative dance generation and editing.
Figure 7: Music-conditioned dance generation of arbitrary Style. In this example, three generated dances are conditioned on the same input music but with different style prompts.
IV-H Applications
This work is applicable to many interesting and promising applications, some of which are outlined below.
Interactive Dance Generation: The proposed method can be applied to interactive dance generation by allowing users to adjust style, timing, and movement constraints, enabling personalized and adaptive dance performances to meet artistic and practical needs.
Creative Dance Generation: Our method has the potential for creative dance generation, which emphasizes the production of movements with artistic style, rhythm, and emotional expression. This task typically involves modeling complex spatial-temporal patterns and promoting motion diversity beyond mere physical realism.
Dance Generation of Arbitrary Style: Our method can generate dances in arbitrary styles based on conditioned music, enabling diverse dance generation. Spatial constraints can also be applied to adjust the generated dance movements to meet user-defined requirements or creative contexts. Visualizations of our approach for dance generation in arbitrary styles are shown in Figure 7.
V Conclusion
@@
This work studies controllable dance generation, an important yet often neglected task. We introduce Style-Guided Motion Diffusion (SGMD), which enhances the existing diffusion-based dance generation framework with a Style Modulation module and Spatial-Temporal Masking. We establish new experimental settings for controllable dance generation under different spatial constraints. Extensive experiments demonstrate that our approach achieves state-of-the-art performance across a variety of tasks, including long-term dance generation, trajectory-based dance generation, dance in-betweening, and dance inpainting. We also explore different style prompts, including one-hot encoding, genre name, and style description, and experimentally find that style description, with rich semantic information, yield the best performance. Ablated experiments demonstrate the effectiveness of the modulation module and style prompts. We also compare results of dance generation using different audio feature extractors. To obtain dances that align better with the music, we suggest using Jukebox to extract audio representations and applying style description prompts augmented by large language models. The proposed framework is suitable for flexible applications of controllable dance generation, enabling the user to freely edit or generate the desired dance sequence. We hope that this work will inspire further research in the fields of automatic and interactive dance generation.